# A Video is Worth 10,000 Words: Training and Benchmarking with Diverse Captions for Better Long Video Retrieval

### v1

@@ 4.1 Method
We propose losses complementary to any model that uses contrastive-style loss, in our case, finetuning both COSA and InternVideo. We illustrate this approach in Figure 4. We sample a batch of videos with corresponding captions and apply a loss that pushes matching video and caption embeddings closer together. To encourage the model to associate all of the descriptions for a video with that video we add the additional 9 captions for a video during training.
Figure 5: We measure the extent to which our data augmentation improves performance in the data constrained regime, training only with the indicated amounts of data, and computing paragraph-to-video and 10k Words retrieval for the resulting trained models. We show, for ActivityNet10k, that our finetuning method (ours) is consistently better than vanilla COSA finetuning (C) for retrieval with Full, Long, and Short captions.
[…]
$L=l_{\text{itc}}+\alpha_{\text{t2t}}\cdot l_{\text{t2t}}+\alpha_{\text{proj}}\cdot l_{\text{proj}}$ (1)
@@
We use 3 settings in this paper: “Domain Finetune” which is just the default setting of whichever model (COSA or InternVideo) we are finetuning, “10k Finetune” where we set $\eta=0.75$ but keep $\alpha_{\text{t2t}},\alpha_{\text{proj}}=0$, and then “Ours” where we set $\eta=0.75$ and $\alpha_{\text{t2t}},\alpha_{\text{proj}}=0.1$. We set these values after some ablations, which we report in the supplement for reference.
Table 10: LFVILA COSA finetuning results.  See Section 4.1 for explanations of settings. 10k Words data helps performance.
Finetune Method All Short Long Partial
R@1 Avg. R R@1 Avg. R R@1 Avg. R R@1 Avg. R
Domain 77.3 86.9 65.2 78.4 90.2 95.9 73.8 84.9
10k 85.2 92.6 78.2 89.2 95.3 98.2 73.0 83.9
Ours 84.9 92.5 78.0 89.1 95.0 98.1 71.9 83.3
Table 11: ActivityNet InternVideo finetuning results. Our method improves performance here as well, although the majority of the improvement here can be attributed to the data.
@@
10k 42.2 63.6 52.6 72.8 30.0 52.6 49.1 70.5
Ours 42.2 63.7 52.6 72.9 30.0 52.7 49.4 70.7
4.2 Results
@@
Table 8 shows the results for COSA finetuning on ActivityNet (”Domain”). We observe that finetuning by sampling from 10k Words data (”10 Words”) yields considerable improvements for retrieving with 10k Words captions. We also see that introducing our novel losses (”Ours”) further improves performance. We show that these findings hold when finetuning on other datasets, such as LF-VILA (Table 10), as well as when we adapt our data sampling and losses for another state-of-the-art model, InternVideo, in Table 11. This also reveals that the LF-VILA dataset is easier to tackle once the model can finetune. This is perhaps an artifact of LF-VILA’s original automatically generated captions, or due to the fact that LF-VILA is open domain and lacks the fine-grained differences present in ActivityNet.
Table 12: Data sampling ablation. We experiment with using 10k Words data consisting of the indicated types: Simplification (l+e, l+i, l+u), Joint (s+e, s+i, s+u), and/or Partial. We observe that even when the model is trained only with the simplification captions, it has generalizes well for types of captions.
Sampled 10k Types ActivityNet
Simplification Joint partial Full All Short Long
59.3 43.55 30.7 56.4
60.1 45.1 33.2 57.0
58.1 43.1 30.8 55.3
58.6 44.3 33.4 55.1
57.7 43.5 32.7 54.3
58.5 44.2 32.8 55.5
Since part of our contribution is a data augmentation strategy, we also evaluate its performance by finetuning with different fractions of the original ActivityNet data in Figure 5. We find that using 10k Words data consistently improves performance for all amounts of training data.
4.3 Ablations
In Table 12, we find that the finetuning is quite sensitive to the type of data sampled, where using partial data while training our losses tends to be detrimental to the performance, regardless of which caption is used. Notice how finetuning even with only simplification samples, which are paragraph length, results in +0.8% improvement for Short R@1 compared to “Domain” finetuning, which suggests the 10k Words data has nice properties for generalization. We perform ablations for $\eta$, $\alpha_{\text{proj}}$, and $\alpha_{\text{t2t}}$ as well (see appendix, and find that most hyperparameter changes are not more significant than variance across runs, except for setting $\eta=0$, which is detrimental to 10k Words performance.
@@ 5 Related Work
Video-Language Models. Recent efforts have started to focus on building a multimodal foundation model that understands not just vision and language, but captures the extra complexities of video and language. These approaches are typically pre-trained in a manner inspired by CLIP [39] and ALIGN [24], using sets of video and text pairs with varying levels of noise [34, 7, 60], where the text is some form of video description, caption or summary. These models are typically pretrained on some task or set of tasks, and then used on downstream tasks either in a zero-shot manner or after finetuning. Early approaches use pre-computed features to represent videos [59]. Typical models learn a shared embedding space between the videos and text [50, 4, 16, 18, 34, 35, 23]. Others process concatenated video and text inputs with cross-modal encoders [44, 30, 43, 14, 15, 62, 49]. Some even use still images or average frame embeddings and achieve quite strong performance [9, 29, 6]. However, as computational resources have scaled, so have the methods, and many current approaches learn to compute features from raw video [7, 46, 12, 31]. Along with this trend models are branching out from contrastive learning to incorporate other learning tasks as well, even generative objectives including captioning [46, 12, 51, 13, 21, 11, 55, 28]. In this paper we focus on the video retrieval task, and show results using VideoCLIP [50], Frozen [7], COSA [12], and InternVideo [46] as a representative set of video-lanuage models.
@@
Long Video Understanding. Videos in the computer vision literature tend to be short – the average length of videos in tentpole datasets [41, 52, 19, 10, 22] is under 30 seconds. Over the years, some have introduced datasets consisting of longer videos [17, 5, 61, 37, 45]. With the introduction of these datasets, larger GPUs, and advancements in vision-text modeling, many researchers have begun proposing methods that either address long video as a first-class interest [6, 45, 40], or at least, are flexible for both long and short videos [31, 11, 13, 46, 12]. In this paper we focus on long video from the perspective of the retrieval task, and we propose a method for data expansion to allow for more thorough development, evaluation, and benchmarking of long video understanding approaches. This is reminiscent of [47], however, unlike their work, we generate novel synthetic captions, and restrict our training and analysis to the more common single query retrieval setting.
@@
Text Summarization. Our work bears some resemblance to efforts in the areas of controllable text summarization and simplification. For these tasks, control tokens dictate how a model simplifies or summarizes text while preserving its meaning. The definition of these tokens is a key differentiator between papers [2] – they can be user-defined [56, 36, 3, 1, 58], or optimized over some data [33, 32, 38, 26, 42]. Additionally, these pipelines often involve a human-in-the-loop at inference time to give keywords [20] and can require heavy labeling [53, 2]. We opt to use our LLM approach to avoid a reliance on control token definition, human-in-the-loop inference, or depedency on strongly annotated data and the biases that come from training on such limited data. As a result, our outputs are somewhat less “controllable” but are, as shown in Section 2, sufficient for our needs.
@@ 6 Conclusion
We showed how the video retrieval problem can be expanded to capture all ways someone might describe a video. We built ActivityNet10k, QuerYD10k, and LF-VILA10k, whose captions refer to videos using a variety of natural language descriptions instead of just one. We showed that SOTA models fail to generalize well to all potentially valid descriptions, and proposed a novel finetuning approach that can retrieve videos effectively from a range of descriptions. We proposed a finetuning approach to solve that more fine-grained retrieval problem more accurately. This work shows how the complexities of video understanding can be measured and addressed using language as a guide. We hope future work will continue to explore the complete spectrum of language that can be used to capture video content.
@@ 7.1 Prompts and Costs
We reduce the cost in terms of input token counts by batching our inputs. For example, we are generating 3 different summarizations per paragraph, but the source paragraph is the same in all 3 cases. So, instead of passing the input once for each level of summarization (3 times total), we pass the input once, and ask for all summarizations to be present in the output, reducing our input tokens by a factor of 3. So, our final prompts are as follows for summarization, simplification, and joint:
@@
You are a helpful writing assistant, with a speciality in summarizing text-based scene descriptions. You will be asked to write 3 summaries of the scene described in the following paragraph, indicated by PARAGRAPH. Do not modify the indicated order of events. Prioritize visual details. Do not hallucinate. Do not describe objects or events that do not appear in the original paragraph. PARAGRAPH: ORIGINAL PARAGRAPH. Label this summary as SUMMARY_1. For this summary, please write 10 words which summarize the scene described by the PARAGRAPH. Do not use more or less than 10 words. Without using more than 10 words, write complete sentences. Label this summary as SUMMARY_4. For this summary, please write 40 words which summarize the scene described by the PARAGRAPH. Do not use more or less than 40 words. Without using more than 40 words, write complete sentences. Label this summary as SUMMARY_7. For this summary, please write 70 words which summarize the scene described by the PARAGRAPH. Do not use more or less than 70 words. Without using more than 70 words, write complete sentences.
@@
You are a helpful writing assistant, with a speciality in simplifying and rewriting descriptions for different age groups and reading levels. You will be asked to write 3 versions of the scene described in the following paragraph, indicated by PARAGRAPH. Do not modify the indicated order of events. Prioritize visual details. Do not hallucinate. Do not describe objects or events that do not appear in the original paragraph. PARAGRAPH: ORIGINAL PARAGRAPH. Label this version as VERSION_primary_school. For this version, rewrite the PARAGRAPH with 70 words to make it suitable for a primary school reading level. Label this version as VERSION_secondary_school. For this version, rewrite the PARAGRAPH with 70 words to make it suitable for a secondary school reading level. Label this version as VERSION_university. For this version, rewrite the PARAGRAPH with 70 words to make it suitable for a university reading level.
@@
You are a helpful writing assistant, with a speciality in summarizing text-based scene descriptions. You also have a speciality in simplifying and rewriting descriptions for different age groups and reading levels. You will be asked to use 10 words to write 3 summaries of the scene described in the following paragraph, indicated by PARAGRAPH. Do not modify the indicated order of events. Prioritize visual details. Do not hallucinate. Do not describe objects or events that do not appear in the original paragraph. PARAGRAPH: ORIGINAL PARAGRAPH. Label this version as VERSION_primary_school. For this version, rewrite the PARAGRAPH with 10 words to make it suitable for a primary school reading level. Do not use more or less than 10 words. Without using more than 10 words, write complete sentences. Label this version as VERSION_secondary_school. For this version, rewrite the PARAGRAPH with 10 words to make it suitable for a secondary school reading level. Do not use more or less than 10 words. Without using more than 10 words, write complete sentences. Label this version as VERSION_university. For this version, rewrite the PARAGRAPH with 10 words to make it suitable for a university reading level. Do not use more or less than 10 words. Without using more than 10 words, write complete sentences.

### v2

@@ 4.1 Training-time Improvements
We propose an approach for leveraging our 10k data that is lightweight and flexible, allowing us to perform finetuning both COSA and InternVideo. We sample a batch of videos with corresponding captions and apply a loss that pushes matching video and caption embeddings closer together. To encourage the model to associate all of the descriptions for a video with that video we also include the synthetic captions for a video during training.
Specifically, for every video we sample (i) the ground truth paragraph, and (ii) a random 10k Words caption. We mix the two sets of captions, taking one caption per video, to yield our primary text features, $f_{t}$, ensuring that a fixed percentage (set by a mixing ratio, $\eta$) are 10k Words captions, and the rest are ground truth. Using these primary text features, we compute standard bi-directional contrastive loss with the video features as in COSA [12] and InternVideo [46]. The advantage of such a simple formulation, is that it is easily reusable across many SOTA video-language models, since most leverage some video-text contrastive loss. This allows us to apply it to both COSA (Table 8) and InternVideo (Table 10).
@@
We use 2 settings in this paper: “Domain Finetune” which is just the default setting of whichever model (COSA or InternVideo) we are finetuning (with no synthetic captions), and “Ours” where we set $\eta=0.75$. We set these values after some ablations, although ultimately these ablations (see Appendix) suggest that models are not sensitive to the exact $\eta$ so long as it is not extremely high or low.
Table 10: ActivityNet InternVideo finetuning. Ours is best.
@@
Ours 42.2 63.6 52.6 72.8 30.0 52.6 49.1 70.5
Figure 2: We plot standard caption retrieval results for each item in ActivityNet, sorted by rank. We also plot the retrieval for a few synthetic caption types, sorted by standard caption retrieval rank. For many samples, synthetic captions yield superior retrievals.
@@
Table 8 shows the results for COSA finetuning on ActivityNet (“Domain”). We observe that finetuning by sampling from 10k Words data (“Ours”) yields considerable improvements for retrieving with 10k Words captions. We show that these findings hold when we adapt our data sampling and losses for another state-of-the-art model, InternVideo, in Table 10, as well as when finetuning on other datasets, such as LF-VILA (see appendix). While the improvements on InternVideo are less dramatic (perhaps because it is closer to the best-case performance, see Section 5, they are still nontrivial, and give evidence for the effectiveness of our data. We find that LF-VILA10k responds more to finetuning, perhaps because it is open-domain and lacks the fine-grained difficulties of ActivityNet.
Table 11: ANet query ensemble retrieval results. We compare standard retrieval on the zero-shot, domain-finetuned, and 10k-finetuned COSA models to retrieval with an ensemble of synthetic captions. The ensemble is very effective for zero-shot.
Finetuning Inference R@1 Avg. R
Zero-shot Standard 34.2 56.2
Ensemble 37.6 58.6
Domain Standard 59.1 77.7
Ensemble 59.2 77.9
Our Standard 60.1 78.7
Ensemble 60.2 78.8
Table 12: We show some example failures, where the target video is retrieved outside the top 10 despite the short description’s uniqueness and specificity. For each failure, we give the short caption (S), the GT caption of the correct video (C), and an incorrec top 10 retrieval (I).
Type Caption
S Man throws ball, goalkeeper blocks.
C The man threw… As the players throw the ball to the goal, the goalkeeper blocked
the ball. The players swim…
I Two teams play… soccer. A goal is scored… with a bicycle kick…
S Two girls play dress up, laughing and drying their faces on a towel.
C Two girls are playing dress … laughing. One girl dries her face on a towel… second
girl dries her face…
I A woman is sitting in a chair. Another woman… starts clipping and filing the other
woman’s nails.
4.2 Inference-time Improvements
The 10k data can improve performance without the need for training at all. It can be used as a form of query expansion to also improve performance at inference time. We already know from Section 3 that overall, standard captions retrieve videos with more recall than 10k captions. However,in Figure 2, we find that the correct retrievals using standard captions are not a superset of the correct retrievals when using synthetic 10k captions. That is, there exist samples where while the standard caption does not retrieve the video well, some 10k caption does. So, we hypothesize that if we aggregate the predictions by attempting the retrieval with both synthetic and standard captions for each sample (instead of only the standard caption), we can improve the quality of the retrieved results.
Our aggregation method is simple. First, we determine which types of 10k captions we will use (see Appendix). We then compute the standard text-video similarity matrix, as well as a separate text-video similarity matrix for each type of 10k caption we choose. We then add these together, giving $50\%$ weight to the standard text-video matrix, and equal weight to the remaining matrices. We report results for performing this sort of query expansion ensemble with COSA zero-shot, domain finetuning, and ours in Table 12.
5 Understanding Failures
To gather insight into whether our dataset is inherently difficult or whether our models should be able to perform better on the 10k Words benchmark, we analyze why models perform much more poorly for short captions than for long captions. The models we finetune are not pretrained on ActivityNet Captions, so there is clearly a domain gap between the training and testing distributions. Our 10k datasets add an additional domain gap for each axis of augmentation. If this domain gap were the primary difficulty introduced by 10k data, then finetuning on the data as we do in our approach would result in similar performance across the different types of 10k descriptions. Instead we see that performance on short captions is much lower than performance on long captions. We investigate this by focusing on two questions. (i) How does the information in a caption affect model performance? (ii) Is each short caption specific enough to uniquely match the corresponding video?
Figure 3: We measure the length and retrieval uniqueness for short caption retrieval, and find that the highest ranks correlate with captions that have lost their unique information.
Figure 4: We measure uniqueness and plausibility for short captions with bad retrievals. We find that most difficult samples tend to be non-unique and have many plausible correct retrievals.
Information Loss. We experiment with three different ways to practically measure the information content of a caption, including an approach based on counting the number of entities relative to the full caption and an approach based on embedding similarities. In Figure 3 we compute information as length of the short caption (measured by raw number of words) divided by the length of the standard caption, which we find provides very similar results to the other approaches (see Appendix) and is simpler. Our results in Figure 3 show that most bad retrievals (high rank) occur when information amount is relatively low, agreeing with the coarse analysis from Table 8. To understand this further we also report relative word counts conditioned on correctness. For samples where both short and standard captions retrieve the matching video at top-1, the short caption has on average 19.1% as many words as the standard caption. When the standard retrieval works and the short fails to retrieve at top-1, this is 15.6%. Clearly, when the synthetic short caption is closer to the length of the original, the retrieval tends to be easier. However, for some samples the retrieval succeeds when the caption is relatively short, indicating there are other relevant factors.
Uniqueness (Automatic). One such factor, uniqueness is the extent to which the information in the short caption overlaps the matching video in ways it does not overlap other videos. We investigate this automatically by measuring how similar sentence embeddings of short captions are to the standard captions of the top-5 actually retrieved videos (not including the matching video). The uniqueness color scale in Figure 3 is darker when these embeddings are close (less unique) and lighter when they are far apart (more unique). Less unique instances tend to appear at higher (worse) ranks. See the appendix for more details on the calculation. We validate this by comparing the average uniqueness for short captions where the model retrieves the correct video at top-1, 0.304, to the uniqueness of the short captions for the samples where the model fails, 0.239. Longer captions more uniquely identify the videos they are supposed to retrieve. More significantly, the figures show many instances where the caption is short but relatively light in color (more unique), suggesting that many of the shorter captions should be able to recall the correct video.
Uniqueness (Manual). We also investigate the uniqueness factor manually. Two authors annotate the top 5 retrievals for 100 failed retrievals (using short captions, where the ground truth video is not in the top 5 for the Domain Finetuned COSA model). Specifically, they compare the short text caption (query), the standard paragraph caption it came from, and the standard paragraph captions of the top 5 retrieved videos. They decide whether the short caption could reasonably describe the same video as the paragraph caption for each of the 5 retrieved videos (whether a retrieval is “plausible”). They then indicate whether the ground truth matching paragraph is a closer match to the short caption than all the top 5 retrieved videos’ paragraphs (whether the short caption is “unique”). They then resolve differences for all 500 “plausible” and 100 “unique” labels.
We plot the results of this analysis in Figure 4, with the x axis indicating how many top 5 retrievals were plausible and the y axis indicating the actual rank of the correct video. Here we see that the rank of the ground truth video is higher when more plausible alternatives are retrieved; the majority of the “bad” retrievals are ambiguous. Surprisingly, sometimes the retrieval is performed successfully even in the cases where the humans considered the text quite ambiguous (5 “plausible” and not “unique”). Most significantly, there exist failed retrievals for short captions which the human labelers did not consider ambiguous, where none of the top-5 retrievals seem “plausible” and the caption is considered a “unique” match. Models could improve their performance on these captions. Table 12 shows some examples of these. Models struggle with fine-grained domains, like the description of a game of water polo (where players “throw” balls), for which the model seems to tend to retrieve videos of soccer (where players “kick” balls) due to the presence of a goalkeeper. Sometimes, it also seems the model is ignoring other key parts of the short caption, such as the presence of the towel in the second example.
@@ 6 Related Work
Video-Language Models. Video-language models build on image-based vision-language models. These approaches are typically pre-trained in a manner inspired by CLIP [39] and ALIGN [24], using sets of video and text pairs with varying levels of noise [35, 7, 60]. These models are typically pretrained on some task or set of tasks, and then used on downstream tasks either in a zero-shot manner or after finetuning. Early approaches use pre-computed features to represent videos [59]. Typical models learn a shared embedding space between the videos and text [50, 4, 16, 18, 35, 34, 23]. Others process concatenated video and text inputs with cross-modal encoders [44, 30, 43, 14, 15, 62, 49]. Some even use still images or average frame embeddings and achieve quite strong performance [9, 29, 6]. However, as computational resources have scaled, so have the methods, and many current approaches learn to compute features from raw video [7, 46, 12, 31]. Along with this trend models are branching out from contrastive learning to incorporate other learning tasks as well, even generative objectives including captioning [46, 12, 51, 13, 21, 11, 55, 28]. In this paper we focus on the video retrieval task, and show results using VideoCLIP [50], Frozen [7], COSA [12], and InternVideo [46] as a representative set of models.
@@
Long Video Understanding. Videos in the computer vision literature tend to be short – the average length of videos in tentpole datasets [41, 52, 19, 10, 22] is under 30 seconds. Over the years, some have introduced datasets consisting of longer videos [17, 5, 61, 37, 45]. With the introduction of these datasets, larger GPUs, and advancements in vision-text modeling, many researchers have begun proposing methods that either address long video as a first-class interest [6, 45, 40], or at least, are flexible for both long and short videos [31, 11, 13, 46, 12]. In this paper we focus on long video in terms of retrieval, and we propose a method for data expansion to enable better training and understanding of long video models. This is reminiscent of [47]; however, we synthesize novel captions, and restrict our training and analysis to the single query retrieval setting.
@@
Text Summarization. Our work bears some resemblance to efforts in the areas of controllable text summarization and simplification. For these tasks, control tokens dictate how a model simplifies or summarizes text while preserving its meaning. The definition of these tokens is a key differentiator between papers [2] – they can be user-defined [56, 36, 3, 1, 58], or optimized over some data [33, 32, 38, 26, 42]. Additionally, these pipelines often involve a human-in-the-loop at inference time to give keywords [20] and can require heavy labeling [53, 2]. We opt to use our LLM approach to avoid a reliance on control token definition, human-in-the-loop, or specially annotated data.
@@ 7 Conclusion
We showed how the data for the video retrieval problem can be expanded to capture a greater variety ways someone might describe a video, instantiated with ActivityNet10k, QuerYD10k, and LF-VILA10k. We showed that SOTA models fail to generalize well to all potentially valid descriptions, and propose fine-tuning and inference-time approaches to mitigate these shortcomings. We also distilled insights on how some SOTA models struggle with short captions. We hope future work further explores the complete spectrum of language that can describe video content.
Acknowledgements. This work was partially supported by NSF CAREER Award (#2238769) to Abhinav Shrivastava. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright annotation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of NSF or the U.S. Government.
@@ 8.1 Prompts and Costs
We reduce the cost in terms of input token counts by batching our inputs. For example, we are generating 3 different summarizations per paragraph, but the source paragraph is the same in all 3 cases. So, instead of passing the input once for each level of summarization (3 times total), we pass the input once, and ask for all summarizations to be present in the output, reducing our input tokens by a factor of 3. We do the same for simplification and joint. So, if we want to generate summarization, simplification, and joint captions for a given ground truth caption, we must make 3 calls to the API (or, if hosted locally, one would have 3 forward passes). Remarkably, the model did not generate a malformed response a single time; in every case, we received each of the 3 requested outputs, properly tagged. It is worth mentioning these could possibly all be batched for a single pass, although at the time of preparing the dataset, the model was less robust under such conditions. If using our strategy for query expansion, discussed in Section 4.2, one would ideally batch all desired axes for a single pass, for the sake of speed.
The resulting costs can be computed in terms of tokens. The summarization prompt is approximately 180 tokens, not including the paragraph. For the 14,926 ActivityNet videos we consider, whose captions are an average of 49.8 words per caption, this means we submitted approximately 3.5 million input tokens for the 3 levels of summarization. Input tokens for the other two axes can be computed similarly. If using certain proprietary models, one must also consider the cost for output tokens, which can be estimated based on the length of the input paragraph compared to the word counts we provide for each dimension in Table 3. So, our final prompts are as follows for summarization, simplification, and joint. Note the use of “primary school” to generate our “elementary” level captions, and “secondary school” to generate “intermediate” captions.
@@
You are a helpful writing assistant, with a speciality in summarizing text-based scene descriptions. You will be asked to write 3 summaries of the scene described in the following paragraph, indicated by PARAGRAPH. Do not modify the indicated order of events. Prioritize visual details. Do not hallucinate. Do not describe objects or events that do not appear in the original paragraph.
[…]
Label this summary as SUMMARY_1. For this summary, please write 10 words which summarize the scene described by the PARAGRAPH. Do not use more or less than 10 words. Without using more than 10 words, write complete sentences.
Label this summary as SUMMARY_4. For this summary, please write 40 words which summarize the scene described by the PARAGRAPH. Do not use more or less than 40 words. Without using more than 40 words, write complete sentences.
Label this summary as SUMMARY_7. For this summary, please write 70 words which summarize the scene described by the PARAGRAPH. Do not use more or less than 70 words. Without using more than 70 words, write complete sentences.
@@
You are a helpful writing assistant, with a speciality in simplifying and rewriting descriptions for different age groups and reading levels. You will be asked to write 3 versions of the scene described in the following paragraph, indicated by PARAGRAPH. Do not modify the indicated order of events. Prioritize visual details. Do not hallucinate. Do not describe objects or events that do not appear in the original paragraph.
[…]
Label this version as VERSION_primary_school. For this version, rewrite the PARAGRAPH with 70 words to make it suitable for a primary school reading level.
Label this version as VERSION_secondary_school. For this version, rewrite the PARAGRAPH with 70 words to make it suitable for a secondary school reading level.
Label this version as VERSION_university. For this version, rewrite the PARAGRAPH with 70 words to make it suitable for a university reading level.
@@
You are a helpful writing assistant, with a speciality in summarizing text-based scene descriptions. You also have a speciality in simplifying and rewriting descriptions for different age groups and reading levels. You will be asked to use 10 words to write 3 summaries of the scene described in the following paragraph, indicated by PARAGRAPH. Do not modify the indicated order of events. Prioritize visual details. Do not hallucinate. Do not describe objects or events that do not appear in the original paragraph.
[…]
Label this version as VERSION_primary_school. For this version, rewrite the PARAGRAPH with 10 words to make it suitable for a primary school reading level. Do not use more or less than 10 words. Without using more than 10 words, write complete sentences.
Label this version as VERSION_secondary_school. For this version, rewrite the PARAGRAPH with 10 words to make it suitable for a secondary school reading level. Do not use more or less than 10 words. Without using more than 10 words, write complete sentences.
Label this version as VERSION_university. For this version, rewrite the PARAGRAPH with 10 words to make it suitable for a university reading level. Do not use more or less than 10 words. Without using more than 10 words, write complete sentences.